# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v3

@@ C.5. Performance improvements
Table 4. HoTPP computation speed improvements in terms of seconds per batch.
@@
RNN autoreg. inference CT-LSTM inference CT-LSTM train ODE inference ODE train
@@
Appendix D Training details
Appendix E Hyperparameters
Dataset-specific training parameters are listed in Table 5.
Table 5. Training hyperparameters.
Dataset Num epochs Max Seq. Len. Label Emb. Size Hidden Size Head hiddens
Transactions 30 1200 256 512 512 $\rightarrow$ 256
MIMIC-IV 30 64 16 64 64
Retweet 30 264 16 64 64
Amazon 60 94 32 64 64
StackOverflow 60 101 32 64 64
Appendix F Datasets Analysis
@@
Table 6. Time step percentiles.
@@
Appendix G Qualitative Analysis of Predictions
@@
Table 7. Example predictions (labels only).
Method Seq. ID Transactions MIMIC-IV Retweet
IFTPP 0 3, 1, 3, 1, 3, 1, 3, 3, 3 10, 27, 23, 22, 1, 27, 3, 28, 26, 25, 23, 30 1, 1
1 6, 3, 6, 6, 6, 6, 6 11, 10 0, 0, 0, 0, 0, 0, 0, 0, 0
2 3, 1, 3, 1, 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
RMTPP 0 3, 1, 3, 1, 3, 1 10, 27, 23, 22, 1, 30, 27, 5, 8, 9, 16, 15 1, 1, 1, 1
1 6, 6, 6, 6, 6, 6 11, 10, 1 0, 0, 0, 0, 0, 0, 0, 0
2 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
NHP 0 1, 1, 1, 1, 1, 3, 1, 1, 1, 1 6 1, 1, 1, 1
1 1, 1, 6, 6, 6, 6, 6 6, 6, 6 1, 1, 1, 1, 1, 1, 1
2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 6, 6 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
ODE 0 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 10, 1, 3, 5, 2, 4, 6, 1 1, 1, 1, 1
1 6, 6, 6, 6, 6 1, 3, 5, 2, 4, 6, 1, 3, 5, 2, 4, 6 0, 0, 0, 0, 0, 0, 0
2 1, 1, 1, 1, 1, 1, 1 2, 4, 6, 1, 3, 5, 2, 4, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
HYPRO 0 3, 1, 16, 3, 12, 1 23, 27, 22, 1, 28, 26, 25, 23 0, 0, 1, 0
1 1, 32, 6, 6, 1, 6 1, 28, 25, 23, 3, 26, 22, 5 0, 0, 0, 1, 1, 0, 1
2 3, 5, 1, 3, 5 15, 13, 2, 7, 4, 1, 3, 5 0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0
Diffusion 0 1, 3, 1, 1, 1, 3, 18, 1 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0
1 6, 6, 6, 6, 6, 6, 6 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 0, 0, 1, 1, 1, 1, 0, 1, 0
2 8, 1, 3, 8, 3, 8, 1, 3, 8, 1, 1, 8 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1
Appendix H T-mAP for Highly Irregular Sequences
Figure 8. Comparison of simple baselines on the Toy dataset with highly irregular time intervals. MAE and OTD metrics represent error values, while T-mAP measures model quality.
Appendix I Long-tail Prediction
Figure 9. Comparison of various metrics on the Transactions dataset across subsets with fewer event classes.

### v4

@@ Appendix G Performance improvements
Table 3: HoTPP computation speed improvements in terms of seconds per batch.
@@
RNN autoreg.
inference
CT-LSTM
inference
CT-LSTM
train
ODE
inference
ODE
train
@@
Appendix H Datasets Analysis
@@
Table 4: Time step percentiles.
@@
Appendix I Training details
Appendix J Hyperparameters
Dataset-specific training parameters are listed in Table 5:
Table 5: Training hyperparameters.
Dataset Num epochs Max Seq. Len. Label Emb. Size Hidden Size Head hiddens Transactions 30 1200 256 512 512 $\rightarrow$ 256 MIMIC-IV 30 64 16 64 64 Retweet 30 264 16 64 64 Amazon 60 94 32 64 64 StackOverflow 60 101 32 64 64
@@ Appendix I Training details
Appendix K Predictions Diversity
Figure 11: Comparison of predicted label entropies, with the dashed horizontal line indicating the ground truth entropy.
@@
The quantitative results for all 5 datasets are presented in Figure 11. For details, please refer to Section 5.3.
Table 6: Example predictions (labels only).
Method Seq. ID Transactions MIMIC-IV Retweet IFTPP 0 3, 1, 3, 1, 3, 1, 3, 3, 3 10, 27, 23, 22, 1, 27, 3, 28, 26, 25, 23, 30 1, 1 1 6, 3, 6, 6, 6, 6, 6 11, 10 0, 0, 0, 0, 0, 0, 0, 0, 0 2 3, 1, 3, 1, 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 RMTPP 0 3, 1, 3, 1, 3, 1 10, 27, 23, 22, 1, 30, 27, 5, 8, 9, 16, 15 1, 1, 1, 1 1 6, 6, 6, 6, 6, 6 11, 10, 1 0, 0, 0, 0, 0, 0, 0, 0 2 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 NHP 0 1, 1, 1, 1, 1, 3, 1, 1, 1, 1 6 1, 1, 1, 1 1 1, 1, 6, 6, 6, 6, 6 6, 6, 6 1, 1, 1, 1, 1, 1, 1 2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 6, 6 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ODE 0 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 10, 1, 3, 5, 2, 4, 6, 1 1, 1, 1, 1 1 6, 6, 6, 6, 6 1, 3, 5, 2, 4, 6, 1, 3, 5, 2, 4, 6 0, 0, 0, 0, 0, 0, 0 2 1, 1, 1, 1, 1, 1, 1 2, 4, 6, 1, 3, 5, 2, 4, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 HYPRO 0 3, 1, 16, 3, 12, 1 23, 27, 22, 1, 28, 26, 25, 23 0, 0, 1, 0 1 1, 32, 6, 6, 1, 6 1, 28, 25, 23, 3, 26, 22, 5 0, 0, 0, 1, 1, 0, 1 2 3, 5, 1, 3, 5 15, 13, 2, 7, 4, 1, 3, 5 0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0 Diffusion 0 1, 3, 1, 1, 1, 3, 18, 1 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1 6, 6, 6, 6, 6, 6, 6 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 0, 0, 1, 1, 1, 1, 0, 1, 0 2 8, 1, 3, 8, 3, 8, 1, 3, 8, 1, 1, 8 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1